Data on the compressive strength of a concrete block after testing are as follows: 2.50 MPa, 2.55 MPa, 2.52 MPa, 2.51 MPa, and 2.56 MPa. Find the absolute uncertainty.


1
Expert's answer
2021-09-28T11:13:05-0400

First, find the deviation of each value from the average value "\mu":


"\mu=\frac{2.50 + 2.55 +2.52 +2.51 + 2.56}{5}=2.53\text{ MPa}."

Find the deviations:

"\delta_1=|\mu-x_1|=2.53-2.50=0.03\text{ MPa},\\
\delta_2=|\mu-x_2|=|2.53-2.55|=0.02\text{ MPa},\\
\delta_3=0.01\text{ MPa},\\
\delta_4=0.02\text{ MPa},\\
\delta_5=0.03\text{ MPa}."

The absolute uncertainty is


"\delta=\frac{\sum^n_{i=1}\delta_i}{n}=\frac{0.03+0.02+0.01+0.02+0.03}{5}=0.02\text{ MPa}."